We have m = 5 preliminary samples of size n = 3 (some numbers have unfortunately
been erased by accident by a clumsy co-op student):We have m = 5 preliminary samples of size n = 3 (some numbers have unfortunately been erased by accident by a clumsy co-op st
Answer #1

Answer * Given that m = 5 n=3 .- Hi, I re;, 2 rei, 3 hi ४ Si I 27.1 29.4.127.2 27.9 12.31 1.3 2 30.6 32.5 32.4 31.83 log 1-07۹) TN ) lel 6.4 1.693 (6 = 3.780 [here, (d2 = 1.693) which is obtained from the control Chart table] .LOL m = 29.0 26 5* 3.78m o UOL - Ť + = 29.026 + 5X 3.780 3 = 29. 026 + 18.9 17320 = 29.026 + 10.9122 UCL - 39.938